solve the equations algebraically:
1/3x+1/3y=1
2x+2y=9

Answer by HyperBrain(694)

The first equation is:
1/3x+1/3y=1
***if we multiply both sides of the equation by 3,
x+y=3
***if we multiply both sides of this equation by 2,
2x+2y=6
++++++++
But, according to the second equation,
2x+2y=9
---this leads to a conclusion that
6=9
which is a false statement.
````
Therefore, there is no solution.